OpenStudy (anonymous):

(4/x-1) + (5)/[(x+1)(x-1)] **** Would i be able to cross out the (x-1) in the denomerator?******

OpenStudy (nikvist):

\[\frac{4}{x-1}+\frac{5}{(x+1)(x-1)}=\frac{1}{x-1}\left(4+\frac{5}{x+1}\right)\]